archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    cut is expected Date Friday 18 March 2011 23 59PM Reason for closing Done Comments 4 Comment by OVH Friday 18 March 2011 23 54PM We will test the cable A break for a few seconds is expected Comment by OVH Friday 18 March 2011 23 55PM Cable is being replaced Comment by OVH Friday 18 March 2011 23 55PM The replacement of the cable has improved the situation but

    Original URL path: http://status.ovh.es/?do=details&id=1244 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    forwarded access difficulties to infrastructures Blizzard After making contact it appears that some of our ranges of ips were indeed blacklisted by Blizzard Concerning the DSL unbundling and collection we now consider that the problem is solved Would you please

    Original URL path: http://status.ovh.es/?do=details&id=1419 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    side The switch is done Comments 11 Comment by OVH Tuesday 31 May 2011 20 43PM Some problems persist on the routing between the collecting operator and OVH The incident is being solved between our teams Comment by OVH Thursday 03 November 2011 13 53PM There are problems persisting on the routing between the collect operator and OVH The incident is being resolved between our teams Comment by OVH Friday 09 December 2011 01 50AM We will make some tests on the links of tests in order to validate the architecture Comment by OVH Friday 09 December 2011 01 52AM We modified the network characteristics of the links between OVH and SFR in order to make our tests The LNS Paris Roubaix Marseille Bordeaux Lyon are ready Comment by OVH Friday 09 December 2011 01 59AM The migration from SFR s side was done Now we are going through the concentration BAS located in Paris The connections are directly established from the DSLAM if it is configured on Level 3 or by BAS as close as possible if the DSLAM are on Level 2 The old sessions are still connected by the BAS in Paris For the next reconnection the

    Original URL path: http://status.ovh.es/?do=details&id=1493 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    ADSL SDSL Task Type Modernization Category Backend Core Status CLOSED Percent Complete Details We are going to update tonight the Dslam lib59 1 The maintenance operation is expected for midnight A 10 minutes interruption is expected Date Tuesday 07 June

    Original URL path: http://status.ovh.es/?do=details&id=1515 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details We are performing maintenance on the card There will be about 3 minutes of interruption Date Monday 11 July 2011 13 36PM Reason for closing Done Comments 1

    Original URL path: http://status.ovh.es/?do=details&id=1603 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    to Project ADSL SDSL Task Type Maintenance Category Backend Core Status CLOSED Percent Complete Details We are updating the concerned equipments 10 minutes of unavailability are expected All will be fixed before 00 30h Date Thursday 21 July 2011 02

    Original URL path: http://status.ovh.es/?do=details&id=1642 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details We will update LNS of roubaix tonight at 00h00 We will currently put in hold all the requests of LNS change towards the LNS of roubaix The operations will be re run later on when the LNS

    Original URL path: http://status.ovh.es/?do=details&id=1718 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    11 The maximum cut is 6 o clock Changes in maintenance will complete this task Date Friday 09 September 2011 01 43AM Reason for closing Done Additional comments about closing A new task has been opened http status ovh co uk do details id 1786 PHPSESSID 7a01e6a51d515c519ac4340f13417347 Comments 2 Comment by OVH Thursday 08 September 2011 08 54AM Actually no sites are cut Optical power measures took a long time

    Original URL path: http://status.ovh.es/?do=details&id=1783 (2015-10-15)
    Open archived version from archive



  •